OpenStudy (anonymous):

solve: |2z - 3| = 4z - 1 and check for extraneous solutions.

OpenStudy (amistre64):

split it into 2 parts: 2z - 3 = 4z-1 and 2z - 3 = -(4z-1)

OpenStudy (amistre64):

weed out any z values that make 4z-1 go negative

OpenStudy (anonymous):

okay, did that. what next?

OpenStudy (amistre64):

what values did you get?

OpenStudy (anonymous):

2z - 3=4z-1 and 2z-3=-4z+1

OpenStudy (amistre64):

solve each one for z using algebraic techniques

OpenStudy (anonymous):

z = -1 and z=1.5

OpenStudy (amistre64):

i think graphically we have something akin to this

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D%7C2x-3%7C+and+y%3D4x-1 almost seen it correctly :)

OpenStudy (amistre64):

4z - 1 > 0 4z > 1 z > .25

OpenStudy (amistre64):

so the negative z can be discarded

OpenStudy (anonymous):

ok, but it says to check for extraneous solutions

OpenStudy (amistre64):

2z - 3=4z-1 2z-3=-4z+1 +3 +3 +3 +3 ----------- ------------ 2z = 4z + 2 2z = -4z + 4 -4z -4z +4z +4z ------------ -------------- -2z = 2 6z = 4 z = -1 z = 2/3

OpenStudy (amistre64):

extraneous solutions are "false" solutions

OpenStudy (amistre64):

when z=-1; we have a false statement, so we discard that as extraneous

OpenStudy (anonymous):

is the 2/3 extraneous too?

OpenStudy (amistre64):

no

OpenStudy (amistre64):

2/3 makes a true statement, therefore it is not a false solution

OpenStudy (anonymous):

ok thank you

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

ok, on this question the equation is: 7|8-3h|=21h-49 the first answer i got was: h=2.5 the second was: 0(zero)h=-1 what does that mean?

OpenStudy (amistre64):

first divide off the 7 7|8-3h|=21h-49 /7 /7 ---------------- |8-3h| = 3h - 7 since |...| are always 0 are bigger; no negative values allowed 3h - 7 >= 0 +7 +7 ------------- 3h >= 7 h >= 7/3 anything less than 7/3 is useless to us

OpenStudy (amistre64):

since |a| = n ; then |-n| or |n| = n a = n or a=-n |8-3h| = 3h - 7 8-3h = 3h - 7 8-3h = -(3h - 7)

OpenStudy (amistre64):

8-3h = 3h - 7 8-3h = -(3h - 7) -8 -8 -8 -8 ------------ ---------------- -3h = 3h -15 -3h =-3h - 1 -3h -3h +3h +3h ------------- -------------- -6h = -15 0 = -1 , simply false h = -15/-6 h = 5/2

OpenStudy (amistre64):

is 5/2 >= 7/3 ??

OpenStudy (amistre64):

since 5*3 >= 7*2 15 >= 14 ; then yes; h=5/2 is correct

OpenStudy (amistre64):

not that i approve of "doing all the work"; but it just seems a better way to explain the process ....

OpenStudy (anonymous):

ok thank you again

OpenStudy (anonymous):

Ok sorry to bother you again but i don't know if i did this one right: 3|5t - 1| +9 < or = to 23 and my answers are: t<or= to 3.54 and t>= -3.14